OpenStudy (anonymous):

a3 + b3 = C3 tell me what numbers can we include to it..?

OpenStudy (anonymous):

its actually \[a ^{3} + b ^{3} = c ^{3}\]

OpenStudy (agent0smith):

This was a futurama joke. Google found this http://www.mathpuzzle.com/twocubes.txt

OpenStudy (amistre64):

zero works ....

Parth (parthkohli):

No non-zero integers...

OpenStudy (anonymous):

This is actually a trick question. It's a variation of Fermat's Last Theorem. It states that for x<2, a^x+b^x=c^x does not exist. Fermat didn't write down his proof, so no one knew how he figured it out. It went unsolved for over 300 years. Not too long ago, a man (I forget his name) finally figured it out. We can now prove that there is not a single set of numbers that you can plug into that equation with all three exponents being the same (with an exponent over two).

OpenStudy (unklerhaukus):

Andrew Wiles

OpenStudy (anonymous):

Side note time! Thanks for the medals guys! I wasn't expecting such a response! I love sharing my random extra math knowledge with people, and this is the first place I've actually gotten a good response to it. Also, if you're interested in this topic, I highly suggest Simon Singh's book, Fermat's Enigma. I'm about halfway through it, and I love it. Then again, I absolutely love Simon Singh in general. Anyway, it's an extremely interesting topic and I really think it's worth learning more about. Thanks for listening to me ramble in my nerdy way for a bit!

Parth (parthkohli):

There are infinite solutions.

OpenStudy (anonymous):

No Parth, there are no solutions. This is a Fermat's last theorem question.

Parth (parthkohli):

@singlesixx do you see the asker mentioning anywhere that they wanted integral solutions?

Parth (parthkohli):

\[0^3 + 0^3 = 0^3\]Booyah.

OpenStudy (anonymous):

no we can't

Parth (parthkohli):

\[1^3 + 0^3 = 1^3\]BOOYAH

OpenStudy (anonymous):

no dude

Parth (parthkohli):

@singlesixx The original poster hasn't explicitly asked us for integral non-zero distinct solutions. That's three things you're assuming :-)

OpenStudy (anonymous):

it can't be small numbers because global tried this....

Parth (parthkohli):

@singlesixx Also, we're talking about \(x > 2\). You wrote \(x<2\)

Parth (parthkohli):

@UnkleRhaukus But we gotta love Andrew Wiles, eh? :-)

OpenStudy (unklerhaukus):

\[(i^{2/3})^3 + (5^{1/3})^3 = (2^{2/3})^3\]

Parth (parthkohli):

lol

OpenStudy (anonymous):

Parth, honey, cut it out. You're just mad because I know what I'm talking about. I did not mention integral solutions. I did make a mistake in saying x<2 instead of x>2, but my point was still clear. 0, 1, and 2 are the exponents that are generally accepted to work. Yes, you can get around that by using 1 or 0 as the base, but nobody really thinks of that as a solution. I know what I'm talking about, alright? This is one of my favorite theorems and I know a lot about it. You're trying to undermine me to make yourself look better and it's childish.

OpenStudy (unklerhaukus):

how old exactly do you think the great PK is? @singlesixx

OpenStudy (anonymous):

I have no way of knowing how old he is. Frankly, it doesn't matter.

Parth (parthkohli):

13

OpenStudy (anonymous):

Okay, why does that matter? I know several 13 year olds that are pretty mature. I hope you aren't using your age as an excuse. I'm only three years older than you.

OpenStudy (anonymous):

mmmmm...

OpenStudy (anonymous):

\(a^n+b^n=c^n\) as no integer solutions if \(n\geq 3\) it is fermat's last theorem, but the case for \(n=3\) is not too hard case \(n=4\) is even easier general proof would require probably 10 years of study to being to understand the structure of the proof i imagine

Parth (parthkohli):

\[\large 0^n + 0^n = 0^n\]

OpenStudy (anonymous):

your question is not wrong

OpenStudy (anonymous):

for example i will give you one solution x=1/{1/2}^(1/3), y=1/{1/2}^1/3 and z=1/{1/4}^1/3 these are real values there are also so many real vaues that can satisfy your equation.

OpenStudy (anonymous):

i mean x=a, y=b and z=c

OpenStudy (anonymous):

i asked numbers dude....

OpenStudy (anonymous):

just use calculator

OpenStudy (anonymous):

I would say that any positive integer can be included in this equation. This is simply say that the sum of two cubes is equal to another cube. Any positive number can have a cube root, so?

OpenStudy (anonymous):

not to bump this but it showed up under 'open questions' and I felt compelled to respond... @singlesixx we're all glad that you have 'extra' math knowledge but you needn't take being proven wrong so personally. Although initially I was going to reward you a medal, reading your immature responses to PK's valid challenges (mostly because of the ambiguity with which you stated Fermat's last theorem) changed my mind. Ironically enough you called his behavior 'childish'... you should avoid coming across as arrogant when this website has several mathematicians far better than both you and I :-p

OpenStudy (anonymous):

still i didn't got it.. :(

OpenStudy (anonymous):

Read this http://en.wikipedia.org/wiki/Fermat%27s_Last_Theorem

OpenStudy (anonymous):

great..

OpenStudy (zzr0ck3r):

why is this bumped?

OpenStudy (anonymous):